OpenStudy (anonymous):

A parachutist's rate during a free fall reaches 264 feet per second. What is this rate in meters per second? At this rate, how many meters will the parachutist fall during 15 seconds of free fall? In your computations, assume that 1 meter is equal to 3.3 feet. Do not round your answers.

OpenStudy (anonymous):

@Directrix @kmeis002 @pooja195 @sleepyjess

OpenStudy (anonymous):

for a medal

OpenStudy (anonymous):

@EclipsedStar

OpenStudy (anonymous):

never mind but thanks anyway :)

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

can i give you a medal

OpenStudy (doc.brown):

1 foot = 12 inches 1 inch = 2.54 cm 100 cm = 1 m\[\frac{264ft}{s}\times\frac{12in}{ft}\times\frac{2.54cm}{in}\times\frac{m}{100cm}\]\[\frac{264\cancel{ft}}{s}\times\frac{12\cancel{in}}{\cancel{ft}}\times\frac{2.54\cancel{cm}}{\cancel{in}}\times\frac{m}{100\cancel{cm}}\]\[=80.4672\frac{m}{s}\] Although your teacher wants 1m = 3.3ft for some reason. \[\frac{264ft}{s}\times\frac{m}{33ft}\]\[\frac{264\cancel{ft}}{s}\times\frac{m}{3.3\cancel{ft}}\]\[=80m/s\]

OpenStudy (doc.brown):

\[\frac{80m}{s}\times\frac{15}{15}=\frac{1200m}{15s}\]Assuming no change in velocity, in 15s the parachutist will have traveled 1200m.